OpenStudy (anonymous):

Which expression can be used to find the area of a circle with a diameter of 60 cm? A. π · 60 B. π · 302 C. π · 602 D. π · 1202 plz help will give medal

OpenStudy (anonymous):

D=60,r=60/2=30 A=pi*r^2 A=pi*(30)^2

OpenStudy (anonymous):

answer is b pi(30)^2

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

got 100% thanx so much

OpenStudy (anonymous):

no problem